OpenStudy (anonymous):

In quadratic equations, how do you convert an equation from general form (y = a(x-h)^2+k) to standard form (y = ax^2+bx+c)?

OpenStudy (anonymous):

Please show work!

OpenStudy (mathstudent55):

Square the binomial. Then distribute the 1/2. Then combine like terms and write it in descending order of degree.

OpenStudy (anonymous):

Where does the x come from in the second monomial? 1/2(x-3)^2+8

OpenStudy (anonymous):

So far I have y = 1/2x^2 - 4.5 + 8? Or is there no x?

OpenStudy (mathstudent55):

\(y = \dfrac{1}{2}(x - 3)^2 + 8\) First, square the binomial x - 3. The square of a binomial has 3 terms and follows the pattern: \( (a + b)^2 = a^2 + 2ab + b^2\) \( (a - b)^2 = a^2 - 2ab + b^2\) You need the second one because you have x - 3, not x + 3.

OpenStudy (mathstudent55):

In general, \((a + b)^2 \ne a^2 + b^2 \) \((a - b)^2 \ne a^2 - b^2 \)

OpenStudy (anonymous):

So that would be 1/2(x^2 - 6x - 9) + 8 Which would be 1/2x^2 - 3x - 4.5 +8 Then you would get 1/2x^2 - 3x + 3.5 Right?

OpenStudy (mathstudent55):

Almost. The sign of the 9 is +9, not -9 when you square the binomial.

OpenStudy (anonymous):

Okay, thanks!

OpenStudy (anonymous):

I understand how to do it now.

OpenStudy (mathstudent55):

\(y = \dfrac{1}{2}(x - 3)^2 + 8\) \(y = \dfrac{1}{2}(x^2 - 6x + 9) + 8 \) \(y = 0.5x^2 - 3x + 4.5 + 8 \) \(y = 0.5x^2 - 3x + 12.5 \)

OpenStudy (mathstudent55):

Great.